# Smart walkers in discrete space

### v1

@@ 1 Introduction
The study of random walkers has historically provided invaluable insights into a wide range of complex phenomena. Examples can be found in statistical physics [18], network theory [12, 7, 11], robotics [23] or finance [4]. Random walkers often navigate in discrete spaces by using simple stochastic rules. Therefore, they offer a reference baseline to elaborate on a large plethora of complex dynamical schemes. These latter may account for adaptability, memory, or even strategic interactions to enable for a comprehensive description of the different phenomena under scrutiny.
@@
The manuscript is structured as follows: In Sec. 2 we describe the model of two random walkers on a Markov chain as a game. In Sec. 3, we present the simplest case of two random walkers. We provide an analytical derivation of the probability distribution of the first encounter as a new result, and then we re-derive known results [13] through a novel and simpler proof detailed in Appendix A. This result constitutes the benchmark reference for the subsequent analysis. In Sec. 4, we illustrate how to incorporate a smart strategy into a walker, outlining the reinforcement learning training procedure. We also introduce the information-theoretic and thermodynamic metrics utilized in our analysis. Sec. 4.2 presents our numerical results, achieved by training a random walker using linear, sinusoidal, and non-Markovian reward policies. Moreover, in Sec. 5, and motivated by the success of the above analysis, we speculate that configuration entropy can be employed as an effective measure of attitude or intelligent behaviour. To elaborate in this direction we also consider a specific application to table top games. Finally, we conclude in Sec. 6 by discussing the limitations and strengths of our research.
@@
(b) First encounter probability distribution after training Alice under linear time-dependent reward signal (4.7).
Figure 6: First encounter probability distributions, derived both numerically (10,00010{,}000 simulated games after learning) and analytically with (3.6).
2 The reference setting
3 The case of two simultaneously evolving random walkers
@@
Starting from these premises, in the next Section we will set to consider the generalized case study where one walker is allowed to recast its own dynamics in light of the gained experience. The smart agent is exposed to different quenched rewards, as we shall highlight in the following, and training is implemented via reinforcement learning algorithms.
@@ 4.1 The training pipeline
where the state–action quality tensor (Q-tensor) is updated using q-learning: a temporal-difference learning rule, with a one-step look-ahead into the future. The parameter $\alpha\in[0,1]$, known as the learning rate, determines the extent to which new information overrides prior knowledge. It governs how aggressively the Q-values are updated based on new experiences. The variable $r$ denotes the reward received by the agent after taking action $a$ in state $s$; in our setup, this corresponds to the score assigned at the resulting meeting position. If no score is gained, the reward is simply set to zero. Finally, the parameter $\gamma\in[0,1]$, known as the discount factor, controls the agent’s preference for future rewards over immediate ones. A value of $\gamma$ close to 1 places more emphasis on long-term returns, while values closer to 0 make the agent more short-sighted. The symbols $s^{\prime}$ and $a^{\prime}$ represent the subsequent state and actions following the transition from $s$ via action $a$. We set $\gamma=1$ and $\alpha=\frac{1}{2}$ in our simulations.
We define the agent’s policy $\pi(a|s)$ as a probability distribution over available actions given a state. This distribution is obtained using the Boltzmann exploration paradigm [3], which applies a weighted softmax function to the Q-values of the given state. The weighted softmax is defined as
@@ 4.2 The imprint of learning on the recorded statistics
The agent was trained for 10,00010{,}000 episodes under each reward function, using an inverse temperature parameter $\beta$ that started at an initial value of $1$ and was capped at an upper threshold of $10$. This training regime yielded dynamical systems of the chaser-target typology with substantially morphed first-encounter probability distributions (see Figure 6) and mean first-encounter times (see Figure 7). In the aforementioned Figures, the histogram refer to actual simulation of the learned stochastic dynamics. The solid (red) lines stand for the analytical solutions, computed following the procedure introduced in Section 3, and by making use of the policy table to which Alice converged upon learning, subject to the different assumed rewards.
@@ 5 Configuration Entropy as a Measure of Skill
Building on the results presented in the previous section, it is thus tempting to interpret configuration entropy as an indirect proxy of the "intelligence" of an agent. Notice that the above interpretation is solely justified for the context here analyzed, where just one agent can learn from experience when facing a stereotyped (or unintelligent) opponent. To further challenge this working hypothesis, we move beyond the simple walkers framework to consider a richer class of environments, more complex and widely studied. These are tabletop games.

### v2

@@ 1 Introduction
The study of random walkers has historically provided invaluable insights into a wide range of complex phenomena. Examples can be found in statistical physics [19], network theory [13, 8, 12], robotics [24] or finance [5, 4]. Random walkers often navigate in discrete spaces by using simple stochastic rules. Therefore, they offer a reference baseline to elaborate on a large plethora of complex dynamical schemes. These latter may account for adaptability, memory, or even strategic interactions to enable for a comprehensive description of the different phenomena under scrutiny.
@@
The manuscript is structured as follows: In Sec. 2 we describe the model of two random walkers on a Markov chain as a game. In Sec. 3, we present the simplest case of two random walkers. We provide an analytical derivation of the probability distribution of the first encounter as a new result, and then we re-derive known results [14] through a novel and simpler proof detailed in Appendix A. This result constitutes the benchmark reference for the subsequent analysis. In Sec. 4, we illustrate how to incorporate a smart strategy into a walker, outlining the reinforcement learning training procedure. We also introduce the information-theoretic and thermodynamic metrics utilized in our analysis. Sec. 4.2 presents our numerical results, achieved by training a random walker using linear, sinusoidal, and non-Markovian reward policies. Moreover, in Sec. 5, and motivated by the success of the above analysis, we speculate that configuration entropy can be employed as an effective measure of attitude or intelligent behaviour. To elaborate in this direction we also consider a specific application to table top games. Finally, we conclude in Sec. 6 by discussing the limitations and strengths of our research. The case of two simultaneously trained agents ( interacting via a predator- prey scheme) is shortly addressed in the annexed Appendix C to show the versatility of the proposed approach, beyond the main setting here addressed.
2 The general model
@@ 3 Mathematical setting: two co-evolving random walkers
Starting from these premises, in the next Section we will set to consider the generalized case study where one walker is allowed to recast its own dynamics in light of the gained experience. The smart agent is exposed to different quenched rewards, as we shall highlight in the following, and training is implemented via reinforcement learning algorithms. For the sake of simplicity, we will primarily focus on contexts with just one trainable agent present. However, the introduced formalism can be straightforwardly extended to settings where multiple learners are present (as demonstrated in Appendix C).
@@ 4 The case of a smart walker
(b) First encounter probability distribution after training Alice under linear time-dependent reward (4.8).
Figure 6: First encounter probability distributions, derived both numerically (10,00010{,}000 simulated games after learning) and analytically with (3.8). Each bin of the bar plots corresponds to a different site in the game space. Stated differently, we do not impose any a posteriori binning; the level of displayed coarse graining reflects the specificity of the examined problem. The red curves superimposed to the blue bar plots refers to the theoretical prediction (3.8), while the bar plot reports on the simulation’s result. The theoretical curve takes values on the very same discrete support that defines the available configuration space. No fitting parameters are adjusted when superposing the aforementioned theoretical and numerical curves.
@@ 4.1 The training pipeline
where the state–action quality tensor (Q-tensor) is updated using q-learning: a temporal-difference learning rule, with a one-step look-ahead into the future. The parameter $\alpha\in[0,1]$, known as the learning rate, determines the extent to which new information overrides prior knowledge. It governs how aggressively the Q-values are updated based on new experiences. The variable $r$ denotes the reward received by the agent after taking action $a$ in state $s$; in our setup, this corresponds to the score assigned at the resulting meeting position. If no score is gained, the reward is simply set to zero. Finally, the parameter $\gamma\in[0,1]$, known as the discount factor, controls the agent’s preference for future rewards over immediate ones. A value of $\gamma$ close to 1 places more emphasis on long-term returns, while values closer to 0 make the agent more short-sighted; this can be also leveraged to promote general rapidity (for example quick encounters in chaser-target games). The symbols $s^{\prime}$ and $a^{\prime}$ represent the subsequent state and actions following the transition from $s$ via action $a$. We set $\gamma=1$ and $\alpha=\frac{1}{2}$ in our simulations.
We define the agent’s policy $\pi(a|s)$ as a probability distribution over available actions given a state. This latter distribution is obtained by means of the so called Boltzmann exploration paradigm [3], which applies a weighted softmax function $S_{\beta}$ to the Q-values of the given state. More concretely, we stipulate:
$\pi(a|s)=S_{\beta}(\bm{q}(s,\bm{a}))_{a}$ (4.2)
where $\bm{q}(s,\bm{a})$ stands for the vector of all the qualities corresponding to all the actions $\bm{a}$ that the model can take in state $s$. The weighted softmax $S_{\beta}$ is defined as
@@ 4.2 The imprint of learning on the recorded statistics
The reward profile (4.9) was chosen as it resembles the natural encounter profile of completely random, untrained, agents. The other two linear profiles were instead selected to sketch a toy zero-sum scenario, with or without time pressure. One could interpret, for example, the two walkers as a buyer-seller couple, moving in price space during trading, as already mentioned in the introductory section. The agent (Alice) was trained for 10,00010{,}000 episodes under each reward function, using an inverse temperature parameter $\beta$ that started at an initial value of $1$ and was capped at an upper threshold of $10$. This training regime yielded dynamical systems of the chaser-target typology with substantially morphed first-encounter probability distributions (see Figure 6) and mean first-encounter times (see Figure 7). In the aforementioned Figures, the histogram refer to actual simulation of the learned stochastic dynamics. The solid (red) lines stand for the analytical solutions, computed following the procedure introduced in Section 3, and by making use of the policy table to which Alice converged upon learning, subject to the different assumed rewards.
@@ 5 Configuration Entropy as a measure of skill
Building on the results presented in the previous section, it is thus tempting to interpret configuration entropy as an indirect proxy of the "intelligence" of an agent. Notice that the above interpretation is solely justified for the context here analyzed, where just one agent (in a discrete state space) can learn from experience when facing a stereotyped (or unintelligent) opponent. To further challenge this working hypothesis, we move beyond the simple walkers framework to consider a richer class of environments, more complex and widely studied. These are tabletop games.
@@ Appendix B Get the smart transition matrix
Appendix C Walkers as predator-prey couples
In the main body of the paper, we focused on scenarios were just one agent undergoes training. As already mentioned above, the formalism here presented applies also to settings were multiple walkers are subjected to simultaneous reinforcement learning. To provide a concrete example, in this Appendix we consider a simplified scenario where two walkers are mutually entangled by predator-prey interaction patterns. Specifically, agents are bound to evolve on the very same one dimensional lattice discrete space as employed in the main body of the paper. Therein, we design a reward function that assigns to Alice (the predator) a score +1+1, when an encounter happens. Bob (the prey) is instead given a score 1-1, when hit by Alice. Predator-prey contexts usually impose a time pressure on the agents, and this is accounted for in the analyzed setting by introducing a discount factor $\gamma$ smaller than $1$ (specifically, in the following, $\gamma=0.8$). We also assume that walkers cannot cross each other, as for the other settings here analyzed. We trained both walkers under the above reward pressure: the resulting (after training) empirical distribution of the first encounter is reported in Figure 10, with the same notation adopted for the similar plots discussed in the main body of the paper. Figure 11, on the other hand, displays the agents’ entropy profiles in this new contexts. Since this time both agents undergo reinforcement learning, as we expect both policies to get optimized away from their initial random walking state, yielding a noticeable drop in the respective policy entropy. Also the configuration entropy of the global system drops during training. It can be hence speculated that the effective drop of this latter quantity provides a sensible measure of the collective intelligent (or non random) behaviour, acquired by experience.
Figure 10: First encounter probability distributions, derived both numerically (10,00010{,}000 simulated games after learning) and analytically via Eq. (3.8) (note that the softmax temperature $1/\beta$ is fixed at 0.1). Specifically the bar plot reports the different empirical encounter probabilities, while the red curve refers to the computed theoretical predictions. As for the other setting analyzed throughout the paper, Alice (the predator) starts on the left side, while Bob (the prey) starts on the right. Both walkers converge to the only rational strategy, given the trivial support on which they are both bound: Bob cannot escape the predator since the game board comes to an end at site 1010, but at least he moves all the way to the right to try to delay the encounter, as much as possible. Alice on the other hand has to travel all the way to the right to get the prey. Beyond the game-theoretical interpretation, we remark that the training, as expected, drastically morphs the first encounter probability distribution of the walkers (see Figure 3).
(a) Configuration entropy during training, calculated via (4.5).
(b) Policy entropy of the predator, Alice (via (4.6)).
(c) Policy entropy of the prey, Bob (via (4.6)).
Figure 11: Different entropies are depicted as measured during the training of Alice (predator) and Bob (prey). Notice that, in this new context where agents are simultaneously trained, both entropies (each associated to one of the agents) decrease as a result of the reinforcement learning process. This is also true for the global configuration entropy of the system, which can be therefore conceptualized as an effective measure of the collective level of acquired awareness.